OpenStudy (wesdg1978):

Is this two chain rule problems together or would you use the product rule on this or what? Find g '(x) if g(x)=(x+1)^2(x^2+1)^-3

OpenStudy (mathstudent55):

Use the derivative of a fraction rule.

OpenStudy (wesdg1978):

This might make me look stupid, but I don't see a fraction. What am I missing?

OpenStudy (anonymous):

first, product rule, then chain rule or the fraction rule that is \[g(x)=f(x) / h(x)\] \[g'(x)= \frac{ f'(x)h(x)-h'(x)f(x) }{ h^2(x) }\] but remember to use the chain rule too for the derivatives, by the way x^-3=1/x^3 xD

OpenStudy (mathstudent55):